# The DESI DR1/DR2 evidence for dynamical dark energy is biased by low-redshift supernovae

### v4

@@ 3.2 Systematics diagnosis
This $\sim 0.043$ mag. discrepancy in $-5a_{B}$ within the single DESY5 compilation appears to emerge suddenly around $z\sim 0.1$ where the Universe is still homogeneous, hence we can safely rule out the local inhomogeneous effects. This sudden transition also cannot originate from homogeneous-scale new physics since the PantheonPlus compilation, containing a larger number of low-$z$ well-calibrated SNe, does not admit any such low-$z$ transition. If a homogeneous-scale transition model could compensate for this $a_{B}$ discrepancy within the DESY5 compilation, the same model would, in turn, jeopardize the original $a_{B}$ consistency of PantheonPlus and reproduce a new $a_{B}$ discrepancy within the PantheonPlus compilation. In addition, the 1D probability density distribution (red dashed line) of low-$z$ intercept $-5a_{B}$ in Fig. 2 is non-Gaussian, unlike the other two cases, suggesting the low-$z$ SNe suffer from large scatters and do not follow a consistent magnitude-distance relation. As a comparison, the PantheonPlus SNe contain even more subsamples (18 samples) but still converge to a Gaussian distribution around the weighted average, indicating their well-corrected SNe have followed an accordant magnitude-distance relation just as the DES-SN does. Therefore, the strange transition and large scatter in the low-$z$ intercept $-5a_{B}$ can only be attributed to systematics (including an $M_{B}$ transition). Regardless of its origin, these low-$z$ SNe can only be used for cosmological analyses after correcting this mismatch intercept.
3.3 Systematics corrections
@@ 4 Conclusions and Discussions
Recent cosmological analyses from DESI DR1 and DR2 have attracted wide interest but also raised significant challenges and hot debate to the standard $\Lambda$CDM. In this letter, we reduce the challenge by locating some systematics at low-redshift supernovae. The multi-sample mixed low-$z$ SN sample in DESY5 compilation, suffering from large scatters, fails to rebuild a consistent intercept of the magnitude-distance relation and admits a $\sim 0.043$ magnitude discrepancy with respect to the DES-SN at higher redshifts, which largely drives the preference for dynamical DE. After the first and second versions of our preprint, the DESI DR2 paper (M. Abdul Karim et al., 2025a) has explicitly tested this observation by directly getting rid of low-$z$ SNe in CMB+DESI DR2+DESY5($z>0.1$), where the preference for dynamical DE is reduced to be $2\sigma$ as shown in the middle panel of Fig. 14. After roughly correcting this discrepancy in the intercept for the low-$z$ SN samples, we find the corrected datasets (Planck+DESI DR1/DR2+DESY5) favor dynamical DE only at $1.5\sigma/1.6\sigma$ levels, not significant at all to call for dynamical DE. Therefore, the success of the standard $\Lambda$CDM has not been overturned.
@@
Although we have identified the systematics involving $a_{B}$ inconsistency in the low-$z$ SN sample within the single DESY5 compilation is largely responsible for the immature claim of dynamical DE, the DESI DR1 BAO data alone indeed shows some deviations from the Planck-$\Lambda$CDM model at higher redshifts, especially for those luminous red galaxies (LRG) between $0.80.6<z<$ 0.8 (A. G. Adame et al., 2025a), which have also been considered recently by Z. Wang et al. (2024); G. Liu et al. (2024) as the possible drives for the preference of dynamical DE. This SN-independent deviation from Planck-$\Lambda$CDM model has been re-enforced recently in the DESI DR2 papers (M. Abdul Karim et al., 2025a, b; K. Lodha et al., 2025; W. Elbers et al., 2025) with an explicit discrepancy in $\Omega_{m}$ between Planck-CMB and DESI-BAO and a clear deviation from Planck-$\Lambda$CDM model for LRG2 BAO measurements. Whether this DESI-Planck tension should be better explained by dynamical DE (with phantom crossing) is optional (Z.-K. Guo et al., 2005; B. Feng et al., 2005, 2006), as it can be equally accounted for by some modified gravity with non-minimal coupling (G. Ye et al., 2024; J. Pan & G. Ye, 2025; W. J. Wolf et al., 2025b), mimicking thawing DE (S.-F. Chen et al., 2024; W. J. Wolf et al., 2025a, 2024) even with phantom crossing (R.-G. Cai et al., 2021; A. Chakraborty et al., 2025; J. Khoury et al., 2025).
@@
We thank Tamara Davis, George Efstathiou, Bin Hu, Yipeng Jing, Cheng Zhao, Gong-Bo Zhao, Jian-Qi Liu, and Yan-Hong Yao for helpful discussions. This work is supported by the National Key Research and Development Program of China Grants No. 2021YFA0718304, No. 2021YFC2203004, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 11821505, No. 12235019, No. 11991052, No. 12447101, No. 12073088, and No. 11947302, the China Manned Space Program with Grant No. CMS-CSST-2025-A01, and the Postdoctoral Fellowship Program of CPSF. We also acknowledge the use of the HPC Cluster of ITP-CAS. {contribution} All authors contributed equally to the manuscript.

### v5

@@ III.2 Systematics diagnosis
This $\sim 0.043$ mag. discrepancy in $-5a_{B}$ within the single DESY5 compilation appears to emerge suddenly around $z\sim 0.1$ where the Universe is still homogeneous, hence we can safely rule out the local inhomogeneous effects. This sudden transition also cannot originate from homogeneous-scale new physics since the PantheonPlus compilation, containing a larger number of low-$z$ well-calibrated SNe, does not admit any such low-$z$ transition. If a homogeneous-scale transition model could compensate for this $a_{B}$ discrepancy within the DESY5 compilation, the same model would, in turn, jeopardize the original $a_{B}$ consistency of PantheonPlus and reproduce a new $a_{B}$ discrepancy within the PantheonPlus compilation. This is the key argument of this paper. To explicitly show this, we illustrate in the bottom panel of Fig. 2 with the fiducial cosmology replaced by the best-fit $w_{0}w_{a}$ model from DESI DR2 Abdul Karim et al. (2025a). In this case, not only the low-$z$ and high-$z$ parts of PantheonPlus sample admit different $-5a_{B}$ values, but also the whole intercept $-5_{B}$ is no longer a constant anymore but slightly tilted for both DESY5 and PantheonPlus compilations. In addition, the 1D probability density distribution (red dashed line) of low-$z$ intercept $-5a_{B}$ in Fig. 2 is non-Gaussian, unlike the other two cases, suggesting the low-$z$ SNe suffer from large scatters and do not follow a consistent magnitude-distance relation. As a comparison, the PantheonPlus SNe contain even more subsamples (18 samples) but still converge to a Gaussian distribution around the weighted average, indicating their well-corrected SNe have followed an accordant magnitude-distance relation just as the DES-SN does. Therefore, the strange transition and large scatter in the low-$z$ intercept $-5a_{B}$ can only be attributed to systematics (including an $M_{B}$ transition). Regardless of its origin, these low-$z$ SNe can only be used for cosmological analyses after correcting this mismatch intercept.
III.3 Systematics corrections
@@
It is worth noting that debiasing the $a_{B}$ fluctuation with (3) for low-$z$ sample would slightly change their correlations to DES-SN, so does the fully debiased low-$z$ sample with (III.3). However, this would not affect our conclusion since debiasing the $a_{B}$ discrepancy with (III.3) alone has already diminish the dark energy significance as shown shortly below, and shifting the apparent magnitudes of low-$z$ SNe by the same amount $0.043$ in (III.3) would not change their correlations to DES-SN.
@@ IV Conclusions and Discussions
Recent cosmological analyses from DESI DR1 and DR2 have attracted wide interest but also raised significant challenges and hot debate to the standard $\Lambda$CDM. In this paper, we reduce the challenge by locating some systematics at low-redshift supernovae. The multi-sample mixed low-$z$ SN sample in DESY5 compilation, suffering from large scatters, fails to rebuild a consistent intercept of the magnitude-distance relation and admits a $\sim 0.043$ magnitude discrepancy with respect to the DES-SN at higher redshifts, which largely drives the preference for dynamical DE. After the first and second versions of our preprint, the DESI DR2 paper Abdul Karim et al. (2025a) has explicitly tested this observation by directly getting rid of low-$z$ SNe in CMB+DESI DR2+DESY5($z>0.1$), where the preference for dynamical DE is reduced to be $2\sigma$ as shown in the middle panel of Fig. 14. After roughly correcting this discrepancy in the intercept for the low-$z$ SN samples, we find the corrected datasets (Planck+DESI DR1/DR2+DESY5) favor dynamical DE only at $1.5\sigma/1.6\sigma$ levels, not significant at all to call for dynamical DE. Therefore, the success of the standard $\Lambda$CDM has not been overturned.
@@
Although we have identified the systematics involving $a_{B}$ inconsistency in the low-$z$ SN sample within the single DESY5 compilation is largely responsible for the immature claim of dynamical DE, the DESI DR1 BAO data alone indeed shows some deviations from the Planck-$\Lambda$CDM model at higher redshifts, especially for those luminous red galaxies (LRG) between $0.80.6<z<$ 0.8 Adame et al. (2025a), which have also been considered recently in Refs. Wang et al. (2024b); Liu et al. (2024) as the possible drives for the preference of dynamical DE. This SN-independent deviation from Planck-$\Lambda$CDM model has been re-enforced recently in the DESI DR2 papers Abdul Karim et al. (2025a, b); Lodha et al. (2025b); Elbers et al. (2025) with an explicit discrepancy in $\Omega_{m}$ between Planck-CMB and DESI-BAO and a clear deviation from Planck-$\Lambda$CDM model for LRG2 BAO measurements. Whether this DESI-Planck tension should be better explained by dynamical DE (with phantom crossing) is optional Guo et al. (2005); Feng et al. (2005, 2006), as it can be equally accounted for by some modified gravity with non-minimal coupling Ye et al. (2025); Pan and Ye (2025); Wolf et al. (2025a), mimicking thawing DE Chen et al. (2024); Wolf et al. (2025b, 2024) even with phantom crossing Cai et al. (2021); Chakraborty et al. (2025); Khoury et al. (2025).
Acknowledgements.
@@
We thank Tamara Davis, George Efstathiou, Bin Hu, Yipeng Jing, Cheng Zhao, Gong-Bo Zhao, Jian-Qi Liu, and Yan-Hong Yao for helpful discussions. This work is supported by the National Key Research and Development Program of China Grants No. 2021YFA0718304, No. 2021YFC2203004, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 11821505, No. 12235019, No. 11991052, No. 12447101, No. 12073088, and No. 11947302, the China Manned Space Program with Grant No. CMS-CSST-2025-A01, and the Postdoctoral Fellowship Program of CPSF. We also acknowledge the use of the HPC Cluster of ITP-CAS.